Find the mass of water, in grams, must be added to a 1 mole of potassium nitrate (MM = 101.1 g/mol) to prepare 15%m/m? 


1
Expert's answer
2021-01-18T01:49:00-0500

Let Mass of Water is " m " gram .


So , % = mass KNO3 × 100 / Total Mass .


15 = 101.1 ×100 / ( 101.1 + m )

101.1 + m = 101.1 ×100 / 15


m = 572.9 g .